# Study of charged Lepton Flavor Violation in electron muon interactions

### v1

@@ 3.2 Electron-target experiment with a muon beam
Now we will focus on the low-energy region, where we conducte Monte Carlo simulations to investigate muon-target processes involving $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$. It is essential to highlight that the masses of muon and electron cannot be disregarded, imposing a lower limit on the incident muon’s beam energy, equivalent to its mass. Additionally, the target energy $E_{cm}=\sqrt{2E_{\mu}m_{e}+{m_{\mu}}^{2}+{m_{e}}^{2}}$ possesses its own lower-energy threshold. Consequently, in our simulations, we vary the $M_{Z^{{}^{\prime}}}$ in three distinct sets for each process, ensuring that the muon energy is scanned as close to the lower energy limit as feasible. The outcomes are shown in Fig.LABEL:fig:resonance and Tab.4. Remarkably, a pronounced resonance in the target energy is observed in proximity to $M_{Z^{{}^{\prime}}}$.
@@
Considering the processes we are studying, taking $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ as an example, the main background process corresponding here is $\mu^{+}+e^{-}\to\mu^{+}+\mu^{-}+\gamma$ kahn2018m3 . Correspondingly, its relative rate is $10^{-8}$, therefore we can conduct background-free experimental estimation. Due to $\sigma\propto(\lambda_{e\mu}\times\lambda_{ll})^{2}$and event rate $R=L\cdot\sigma=\frac{dN}{dt}\cdot n_{2}\cdot dx\cdot\sigma$, coupling limits estimates can be made based on the reaction cross-section. As a rough estimate, assuming a 10 cm thick lead target, the incidence rate is about $\frac{dN}{dt}\sim 10^{6}$, the electron number density of lead is about $n_{2}\sim 10^{24}$, and $1\mathrm{y}\sim 10^{7}$\mathrm{s}$$. From this it can be obtained 90% C.L. exclusion lines on the couplings $\lambda_{ee}$ and $\lambda_{\mu\mu}$ products the diagonal coupling $\lambda_{e\mu}$.
The current and prospective limits from low-energy experiments are converted to the coupling limits on $\lambda_{e\mu}\times\lambda_{SM}$ to compare with our results. As shown in Fig.LABEL:fig:3.2.1, in comparison to current results and prospect MEG\@slowromancapii@ experiment, our $ee$ channel simulation can give more stringent upper limits of coupling $\lambda_{\mu e}$. Furthermore, as mentioned in Sec.3.1.2. we also can set new limits on coupling $\lambda_{\mu\mu}\times\lambda_{\mu e}$ as shown in Fig.LABEL:fig:3.2.2, which have never been obtained from experiments yet.
@@ 4 Conclusion
With the continuous development of muon technology, in addition to building high-energy muon collider, there is also certain research prospect for the $e\mu$ interaction. In this work, we investigated the cLFV process propagated by massive neutral gauge bosons ($Z^{\prime}$) in $e\mu$ collision and electron-target experiment with a muon beam, in order to explore the potential of $e\mu$ interactions in new physics searches. We conduct the simulation studies on the cLFV processes $\mu^{+}e^{-}\to e^{+}e^{-}$, $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$, using MadGraph5_aMC@NLO, Pythia8 and Delphes. Then we provide the coupling strength $\lambda_{e\mu}$ and $\lambda_{e\tau}$ at 90% C.L. of different $M_{Z^{\prime}}$. By comparing the sensitivity results with current and prospective limits, it is shown that $e\mu$ interactions have certain research advantages in $\tau$ channel for the heavy $Z^{\prime}$. Furthermore, through the direct interaction of leptons, a unique $\lambda_{e\mu}\times\lambda_{\mu}$ process can be measured, which is unmatched by most other experiments.
@@ Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grant No. 12150005, 12075004, 12175321, 12061141003; the National Key Research and Development Program of China under Grant No. 2018YFA0403900; National College Students Innovation and Entrepreneurship Training Program, Sun Yat-sen University; State Key Laboratory of Nuclear Physics and Technology, Peking University under Grant No. NPT2020KFY04, NPT2020KFY05.

### v3

@@ 3.2 Electron-target experiment with a muon beam
Now we will focus on the low-energy region, where we conduct Monte Carlo simulations to investigate cLFV processes involving $\mu^{+}e^{-}\to e^{+}e^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ in the electron-target experiment with a muon beam. It is essential to highlight that the masses of muon and electron cannot be disregarded, imposing a lower limit on the incident muon beam energy, equivalent to its mass. Additionally, the target experiment energy $E_{cm}=\sqrt{2E_{\mu}m_{e}+{m_{\mu}}^{2}+{m_{e}}^{2}}$ possesses its own lower-energy threshold. Consequently, in our simulations, we vary the $M_{Z^{{}^{\prime}}}$ in three distinct sets for each process as shown in Tab. 4, ensuring that the muon energy is scanned as close to the lower energy limit as feasible. The outcomes are shown in Fig. LABEL:fig:resonance. Remarkably, a pronounced resonance in the target energy is observed in proximity to $M_{Z^{{}^{\prime}}}$.
@@
In this scenario, we conduct a background-free experimental estimation. Due to $\sigma\propto(\lambda_{e\mu}\times\lambda_{ll})^{2}$ and event rate $R=\frac{dN}{dt}\cdot n\cdot dx\cdot\sigma$, where $\frac{dN}{dt}$ denotes the muon production rate, $n$ denotes the electron number density of the target material, $dx$ denotes the thickness of the target, and $\sigma$ denotes the reaction cross section, coupling limits estimates can be made based on the reaction cross-section. As a rough estimate, assuming a 10 cm thick lead target, the incidence rate is about $\frac{dN}{dt}\sim 10^{6}\rm{s^{-1}}$, the electron number density of lead is about $n\sim 10^{24}\rm{cm^{-3}}$, and $1\mathrm{y}\sim 10^{7}$\mathrm{s}$$. From this estimation, we can obtain the 90% C.L. exclusion lines on the couplings $\lambda_{ee}$ and $\lambda_{\mu\mu}$ products the diagonal coupling $\lambda_{e\mu}$.
Similarly, the current and prospective limits from low-energy experiments are converted to the coupling limits on $\lambda_{e\mu}\times\lambda_{\rm SM}$ to compare with our results. As shown in Fig. LABEL:fig:3.2.1, compared with the existing experimental results and expected experimental results, the restrictions given here do not have obvious advantages. However, considering the $\mu\mu$ channel, if the coupling between $Z^{\prime}$ and $\mu\mu$ is much stronger than $e\mu$, that is, $\lambda_{e\mu}\ll\lambda_{\mu\mu}$, and assuming that the coupling strengths of other channels are all set to 0, then $Z^{\prime}$ decay mostly into $\mu\mu$, and the production cross section depends only on $\lambda_{e\mu}$. Accordingly, the resulted limit on $\lambda_{e\mu}$, compared with $\mu\to e\gamma$, can gain advantage of several orders of magnitudes. At the same time, if the data accumulation time of the muon beam experiment is considered to be 10 years, our estimated results will drop by another order of magnitude, which will improve the simulation results to a certain extent. However, it is worth noting that under the above estimation of the coupling strengths, $\Gamma_{Z}$ will be reduced, which means that $Z^{\prime}$ will not be on resonance easily.
Additionally, since the result of $\mu\to e\gamma$ is related to the combination of $m_{e}\lambda_{e\mu}\lambda_{ee}$ and $m_{\mu}\lambda_{\mu\mu}\lambda_{e\mu}$ as Eq. 2, their inference may be non-ignorable. While our detection is exclusively sensitive to the $m_{e}\lambda_{e\mu}\lambda_{ee}$ term or the $m_{\mu}\lambda_{\mu\mu}\lambda_{e\mu}$ term, depending on the channel we choose.
@@ 4 Conclusion
With the continuous development of muon technology, in addition to building high-energy muon collider, there is also certain research prospect for the $e\mu$ interaction. In this work, we investigate the cLFV processes propagated by a massive neutral gauge boson ($Z^{\prime}$) in $e\mu$ collision and electron-target experiment with a muon beam, in order to explore the potential of $e\mu$ interactions in new physics searches. Using the MadGraph5_aMC@NLO, Pythia8 and Delphes software, we conduct the simulation studies on the cLFV processes $\mu^{+}e^{-}\to e^{+}e^{-}$, $\mu^{+}e^{-}\to\mu^{+}\mu^{-}$ and $\mu^{+}e^{-}\to\mu^{+}\tau^{-}$. We then provide the coupling strength $\lambda_{e\mu}$ and $\lambda_{e\tau}$ at the 90% C.L. for different mass of ${Z^{\prime}}$. By comparing the sensitivity results with the limits from the current and prospective experiments, it is shown that $e\mu$ interactions have certain research advantages in $\tau$ channel for the heavy $Z^{\prime}$. Furthermore, based on the universal assumption, a unique process corresponding to coupling $\lambda_{e\mu}\times\lambda_{\mu\mu}$ can be measured, while this coupling combination can only appear in $\mu\to e\gamma$ channel among the current experiments.
@@ Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grant No. 12150005, 12075004, 12175321, 12061141003; the National Key Research and Development Program of China under Grant No. 2018YFA0403900; National College Students Innovation and Entrepreneurship Training Program, Sun Yat-sen University.